ASSIGNMENT OF THE LOWEST EXCITED STATES OF C_{70} AND EVIDENCE FOR FLUORESCENCE FROM THE SINGLET S_{2} STATE

\underline{A. SASSARA}, G. ZERZA and M. CHERGUI, Institut de Physique Expérimentale, Université de Lausanne, CH-1015 Lausanne-Dorigny, Switzerland.

Highly structured fluorescence, phosphorescence and fluorescence-excitation spectra of C_{70} in neon matrices are reported. Fluorescence from both S_{1} and S_{2} states is observed and these are identified as being of A′_{2} and E′_{1} symmetry and have gas phase energies estimated at 15560 and 15725 cm^{−1}, respectively. A third singlet state is identified in the excitation spectrum and is proposed to be of A′′_{2} symmetry with an estimated gas phase energy of 16000 cm^{−1}. Finally, the phosphorescence is dominated by vibrational bands of e′_{1} symmetry pointing to a lowest triplet state of A′_{2} symmetry. Its energy lies at 12588 cm^{−1}above that of the ground state in the neon matrix. From a comparison with data from the literature, the existence of a second triplet state of E′_{1} character is inferred, which lies about 165 cm^{−1}above the lowest triplet state, i.e. with the same energy separation as the lowest two singlet states.